Math 550 (Section 1)  University of South Carolina
D. Meade  Fall 1997

Day One Questionnaire – Math 550 (Fall 1997)

• Personal Information Name Phone Number E-mail Address Major Year

• \underline{Schedule}

Work (hrs/wk) Relax (hrs/wk) Study (hrs/wk) Math 550 (hrs/wk) Other Courses

• Mathematical Background

Course  When  Where  Who  Grade

Math 142 (or equiv)

Math 241 (or equiv)

Other Post-Calculus Math • Other Information

1. Why are you taking this course?

2. Do you have any specific goals or objectives that you hope to fulfill by taking this course?

3. What is your biggest concern about this course?

4. Are you interested in seeing computer demonstrations and learning how to use modern technology (Maple, WWW, Java) to better understand the material in this course?

5. Do you plan to utilize office hours? If “yes”, what are the best times I could have office hours? I’ll try to accommodate as many people as possible, but it is almost impossible to satisfy all requests.